Sheridan Company, organized in 2019, has set up a single account for all intangible assets. The following summary discloses the debit entries that have been recorded during 2020.

1/2/20 Purchased patent (8-year life)

$357,600

4/1/20 Purchase goodwill (indefinite life)

357,000

7/1/20 Purchased franchise with 10-year life; expiration date 7/1/30

437,000

8/1/20 Payment of copyright (5-year life)

164,400

9/1/20 Research and development costs

227,000

$1,543,000

Make the entry as of December 31, 2020, recording any necessary amortization

Reflect all balances as of decemeber 31 2020 using straight line ammortzation
Answer #1
1
Debit Credit
Amortization Expense 80250
      Patents 44700 =357600/8
      Franchise 21850 =437000/10*6/12
      Copyright 13700 =164400/5*5/12
To record amortization
2
Balance of Intangible Assets:
Patents 312900 =357600-44700
Goodwill 357000
Franchise 415150 =437000-21850
Copyright 150700 =164400-13700